# A Review of Modern Recommender Systems Using Generative Models (Gen-RecSys)

### v1

@@ 3.2. LLM-based Generative Recommendation
In LLM-driven generative recommendation, tasks are expressed as token sequences – called prompts – which form an input to an encoder-decoder LLM (e.g. GPT3 (Brown et al., 2020), T5 (Raffel et al., 2020), PALM (Chowdhery et al., 2022)). The LLM then generates another token sequence to address the task – with example outputs including: a recommended list of item titles/ids (Mao et al., 2023; Harte et al., 2023; Sanner et al., 2023; Sileo et al., 2022), a rating (Bao et al., 2023; Kang et al., 2023), or an explanation (Ni et al., 2019; Li et al., 2020; Hada and Shevade, 2021; Geng et al., 2022; Li et al., 2023a). Generative recommendation with LLMs relies on pretraining on large text corpora to provide knowledge about a wide range of entities, human preferences, and commonsense knowledge that can be used directly for recommendation or leveraged to improve generalization and reduce domain-specific data requirements for fine-tuning or prompting (Wei et al., 2022; Bubeck et al., 2023).
@@ 3.2.2. Tuning LLMs for Generative Recommendation
To improve an LLM’s generative recommendation performance and add knowledge to its internal parameters, multiple works focus on fine-tuning (Geng et al., 2022; Bao et al., 2023; Harte et al., 2023; Mao et al., 2023; Kang et al., 2023) and prompt-tuning (Li et al., 2023a; Cui et al., 2022; Zhang et al., 2023b) strategies. Recent works fine-tune LLMs on NL input/output examples constructed from user interaction history and task descriptions for rating prediction (Bao et al., 2023; Kang et al., 2023) and sequential recommendation (Mao et al., 2023; Harte et al., 2023), or in the case of P5 (Geng et al., 2022), both preceding tasks plus top-$k$ recommendation, explanation generation, and review summarization. Other recommendation works study prompt tuning approaches (Li et al., 2023a; Cui et al., 2022; Zhang et al., 2023b), which adjust LLM behaviour by tuning a set of continuous (or soft) prompt vectors as an alternative to tuning LLM weights.
@@
A line of recent work focuses on explanation generation, where prompts are typically constructed from a user’s item or review histories while training explanations are extracted from reviews. Techniques include fine-tuning (Geng et al., 2022; Li et al., 2023a), prompt-tuning (Li et al., 2023a, 2020), and controllable decoding (Ni et al., 2019; Hada and Shevade, 2021), where additional predicted parameters such as ratings steer LLM decoding.
@@ 3.3. Retrieval Augmented Recommendation
While adding knowledge to an LLM internal memory through tuning often improves performance, this approach requires a large number of LLM parameters to store knowledge, and re-tuning is needed for every system update. An alternative paradigm is retrieval-augmented generation (RAG) (Lewis et al., 2020; Izacard and Grave, 2020; Borgeaud et al., 2022), where output generation is conditioned on information retrieved from an external knowledge source such as a dense retriever (Section 3.1). RAG methods can better facilitate online updates, reduce hallucinations, and generally require fewer LLM parameters since knowledge can be externalized (Mialon et al., 2023; Borgeaud et al., 2022; Izacard and Grave, 2020). RAG has recently begun to be explored for recommendation, with the most common approach being to first use a retriever or RS to construct a candidate item set based on a user query or interaction history, and then prompt an encoder-decoder LLM to rerank the candidate set (Hou et al., 2023; Chen, 2023; Wang and Lim, 2023; Dai et al., 2023).
3.4. LLM-based Input Generation
Conversely to how RS or retrievers are used in RAG to obtain inputs for LLMs (Section 3.3), LLMs can also be used to generate inputs for RS (Harte et al., 2023; He et al., 2022b; Rajput et al., 2024; Yuan et al., 2023; Mysore et al., 2023; Li et al., 2023c). For instance: LLM2-BERT4Rec (Harte et al., 2023) initializes BERT4Rec (Section 2.1.1) item embeddings using OpenAI embeddings of item texts; Query-SeqRec (He et al., 2022b) includes LLM query embeddings as inputs to a transformer-based recommender; and TIGER (Rajput et al., 2024) first uses an LLM to embed item text, then quantizes this embedding into a semantic ID, and finally trains a T5-based RS to generate new IDs given a user’s item ID history. Similarly, MINT (Mysore et al., 2023) and GPT4Rec (Li et al., 2023c) produce inputs for a dense retriever by prompting an LLM to generate a query given a user’s interaction history.
@@ 3.5. Conversational Recommendation
The recent advances in LLMs have made fully NL system-user dialogues a feasible and novel recommendation interface, bringing in a new stage of conversational recommendation (ConvRec) research. This direction studies the application of LLMs in multi-turn, multi-task, and mixed-initiative NL recommendation conversations (Friedman et al., 2023; Jannach et al., 2020). It introduces dialogue history as a rich new form of interaction data. Specifically, ConvRec includes the study and integration of diverse conversational elements such as dialogue management, recommendation, explanation, QA, critiquing, and preference elicitation (Jannach et al., 2020; Lyu et al., 2021). While some research (He et al., 2023) approaches ConvRec with a monolithic LLM such as GPT4, other works rely on an LLM to facilitate NL dialogue and integrate calls to a recommender module which generates item recommendations based on dialogue or interaction history (Li et al., 2018a; Chen et al., 2019b; Wang et al., 2022b; Yang et al., 2022). Further research makes progress towards ConvRec system architectures consisting of multiple tool-augmented LLM modules, including components for dialogue management, explanation generation, and retrieval (Friedman et al., 2023; Gao et al., 2023; Wang et al., 2023a, c).
@@ 4. Generative Multimodal Recommendation Systems
In recent years, users have come to expect richer levels of interactions than simple text or image queries, such as providing a picture of their desired product along with a natural language modification (e.g., a dress like the one in the picture but in red). Moreover, users want to better understand the recommendations they receive by visualizing how the product fits their use case (e.g., with a representation of how a garment might look on them, or how a furniture item might look in their room). Such levels of interaction require new RS that can discover the unique attributes or aspects hidden in each modality.
In this section, we discuss RS that utilize multiple data modalities. In Sections 4.1-4.2 we discuss motivations and challenges to the design of multimodal RS. In Sections 4.3-4.4 we review contrastive and generative approaches to multimodal RS, respectively.
@@ 4.1. Why Multimodal Recommendation?
A third use case where multimodal understanding becomes crucial is when considering RS with complex outputs, like those featuring virtual try-on capabilities, or intelligent multimodal conversational shopping assistants.
@@ 4.2. Challenges to Multimodal Recommendation
The development of multimodal RS faces several challenges. First, collecting data to train multimodal systems (e.g., image-text-image triplets) is significantly harder than for unimodal systems. [^3] For example, data augmentation techniques used to define positive pairs for contrastive learning in the unimodal case do not apply to the the multimodal case. As a result, annotations for some modalities may be incomplete (Rahate et al., 2022). Second, combining different data modalities to improve recommendation results is not simple. For instance, existing contrastive learning approaches (Radford et al., 2021; Jia et al., 2021; Li et al., 2021, 2022a) map each data modality to a common latent space in which all modalities are approximately aligned. However, such approaches often capture information that is shared across modalities (e.g., text describing visual attributes), but they overlook complementary aspects that could benefit recommendations (e.g., text describing non visual attributes) (Guo et al., 2019). In general we would like the modalities to compensate for one another and result in a more complete joint representation. While fusion-based approaches (Li et al., 2021, 2022a) do learn a joint multimodal representation, ensuring the alignment of information that is shared and leaving some flexibility to capture complementary information across modalities remains a challenge. Third, learning multimodal models requires orders of magnitude more data than learning models for individual data modalities.
@@ 4.4. Generative Multimodal Recommendation
Diffusion models, explained in Section 2.4, are state-of-the-art models for image generation. While they can also be used for text generation, e.g., by using a discrete latent space with categorical transition probabilities (Austin et al., 2021), text encoders based on transformers or other sequence-to-sequence models are preferred in practice. As a consequence, multimodal models for both text and images, such as text-to-image generation models, combine text encoders with diffusion models for images. For instance, DALL-E (Ramesh et al., 2022) uses the CLIP embedding space as a starting point to generate novel images, and Stable Diffusion (Rombach et al., 2022) uses a UNet autoencoder separately pre-trained on a perceptual loss and a patch-based adversarial objective. Several works have built on and expanded diffusion models by increasing controllability of the generated results (Zhang et al., 2023c), consistency on the generated subjects identity (Ruiz et al., 2023), or for virtual try on applications (Zhu et al., 2023).
@@ 5.1. Evaluating for Offline Impact
The typical approach to evaluating a model involves understanding its accuracy and efficiency in an offline setting, followed by live experiments.
@@ 5.1.1. Accuracy Metrics
The usual metrics used for discriminative tasks are recall@k, precision@k, NDCG@k, AUC, ROC, RMSE, MAE, etc. Many recent works on generative recommender models (e.g., (Kang et al., 2017, 2023; Rajput et al., 2024; Bao et al., 2023; He et al., 2023) incorporate such metrics for discriminative tasks.
@@ 5.1.3. Benchmarks
Many existing benchmark datasets popular in discriminative recommender models, such as Movielens (Harper and Konstan, 2015), Amazon Reviews (He and McAuley, 2016), Yelp Challenge(yel, 2014), Last.fm (Schedl, 2016), and Book-Crossing(Ziegler et al., 2005), are still useful in generative recommender models, but only narrowly. Some recent ones, like ReDial(Li et al., 2018b) and INSPIRED(Hayati et al., 2020), are useful datasets for conversational recommendations. (Deldjoo and Di Noia, 2024; Zhang et al., 2023a) propose benchmarks called cFairLLM and FaiRLLM, to evaluate consumer fairness in LLMs based on the sensitivity of pretrained LLMs to protected attributes in tailoring recommendations. We note that some benchmarks such as BigBench(bench authors, 2023) which are commonly used by the LLM community, have recommendations tasks. We believe it will be specifically useful for the recommendation community to develop new benchmarks for tasks unlocked by generative recommender models.
@@ 5.2. Online and Longitudinal Evaluations
Offline experiments may not capture an accurate picture because of the interdependence of the different models used in the system and other factors. So, A/B experiments help understand the model’s performance along several axes in real-world settings. Note that (Wang et al., 2023e) proposes a new paradigm of using simulation using agents to evaluate recommender models.
In addition to the short-term impact on engagement and satisfaction, the platform owners will be interested in understanding the long-term impact. This can be measured using business metrics such as revenue and engagement (time spent, conversions, etc.). At the same time, several metrics could be used to capture the impact on users (daily/monthly active users, user sentiment, safety, harm, etc.).
@@ 5.3. Conversational Evaluation
BLEU and perplexity are helpful metrics for conversational evaluation and need to be supplanted by task-specific metrics (such as recall) or objective-specific metrics (such as measuring diversity in responses(Li et al., 2015)). Strong LLMs could act as judges, but usually human evaluation is the ultimate choice here. Toolkits such as CRSLab(Kun Zhou, 2021) make it easy to build and evaluate conversational models. However, in many industrial use cases, the lack of labeled data makes this process challenging. Some works  (Friedman et al., 2023) use LLM-powered user simulations to generate data.
5.4. Evaluating for Harm
Milano et al. (2020) proposes six categories of harms associated with RS: content; privacy violations and data misuse; threats to human autonomy and well-being; transparency and accountability; harmful social effects such as filter bubbles, polarisation, and manipulability; and fairness. In addition to these, RS based on generative models may present new challenges (Weidinger et al., 2022; Bird et al., 2023):
  • LLMs use out-of-domain knowledge, introducing different sources of bias that are not easily captured by existing evaluation techniques (Shen et al., 2023; Deldjoo, 2024).
  • Heightened environmental impacts (Berthelot et al., 2024).
@@ 5.5. Holistic Evaluations
As mentioned above, thoroughly evaluating RS for offline metrics, online performance, and harm is highly non-trivial. Moreover, different stakeholders (e.g. platform owners and users)(Abdollahpouri et al., 2020; Sürer et al., 2018; Milano et al., 2021) may approach evaluation differently.
Given the complexity of evaluating Gen-RecSys, there is a significant opportunity for additional research and the development of specialized tooling. Moreover, it would benefit the community to draw inspiration from the HELM benchmark (Liang et al., 2022) and design a similar comprehensive evaluation framework tailored explicitly for Gen-RecSys.
@@ 6. Conclusions and Future Directions
  • Red-teaming – in addition to the standard evaluations, real-world generative RS will have to undergo red-teaming (i.e., adversarial attacks) (Deng et al., 2023; Wang et al., 2023f; shu et al., 2024) before deployment to stress test the system for prompt injections, robustness, alignment verification, and other factors.
Despite being a short survey, in this work, we have attempted to provide a deeper understanding and exploration of the rich landscape of generative models within the field of recommendation systems. This survey extends the discussion beyond large language models (LLMs) to a broad spectrum of generative models and elucidates the nuanced intricacies of their applications across user-item interactions, textual data, and multimodal contexts. Additionally, it illuminates key evaluation challenges and dimensions, addressing the complexities of evaluating Gen-RecSys with considerations for performance, fairness, privacy, and societal impact, thereby establishing a new benchmark for future research in the domain.

### v2

@@ 3.2. LLM-based Generative Recommendation
In LLM-based generative recommendation, tasks are expressed as token sequences – called prompts – which form an input to a seq2seq LLM. The LLM then generates another token sequence to address the task – with example outputs including: a recommended list of item titles/ids (Mao et al., 2023; Harte et al., 2023; Sanner et al., 2023; Sileo et al., 2022), a rating (Bao et al., 2023; Kang et al., 2023), or an explanation (Ni et al., 2019; Li et al., 2020; Hada and Shevade, 2021; Geng et al., 2022; Li et al., 2023a). These methods rely on the pretraining of LLMs on large text corpora to provide knowledge about a wide range of entities, human preferences, and commonsense reasoning that can be used directly for recommendation or leveraged to improve generalization and reduce domain-specific data requirements for fine-tuning or prompting (Wei et al., 2022; Bubeck et al., 2023).
@@ 3.2.2. Tuning LLMs for Generative Recommendation
To improve an LLM’s generative recommendation performance and add knowledge to its internal parameters, multiple works focus on fine-tuning (Geng et al., 2022; Bao et al., 2023; Harte et al., 2023; Mao et al., 2023; Kang et al., 2023) and prompt-tuning (Li et al., 2023a; Cui et al., 2022; Zhang et al., 2023b) strategies. Recent works fine-tune LLMs on NL input/output examples constructed from user-system interaction history and task descriptions for rating prediction (Bao et al., 2023; Kang et al., 2023) and sequential recommendation (Mao et al., 2023; Harte et al., 2023), or in the case of P5 (Geng et al., 2022), both preceding tasks plus top-$k$ recommendation, explanation generation, and review summarization. Other recommendation works study prompt tuning approaches (Li et al., 2023a; Cui et al., 2022; Zhang et al., 2023b), which adjust LLM behaviour by tuning a set of continuous (or soft) prompt vectors as an alternative to tuning internal LLM weights.
@@
A line of recent work focuses on explanation generation where training explanations are extracted from reviews, since reviews often express reasons why a user decided to interact with an item. Techniques include fine-tuning (Geng et al., 2022; Li et al., 2023a; Wang et al., 2024), prompt-tuning (Li et al., 2023a, 2020), chain-of-thought prompting (Rahdari et al., 2024), and controllable decoding (Ni and McAuley, 2018; Ni et al., 2019; Hada and Shevade, 2021; Xie et al., 2023) – where additional predicted parameters such as ratings steer LLM decoding.
@@ 3.3. Retrieval Augmented Recommendation
Adding knowledge to an LLM internal memory through tuning can improve performance, but it requires many parameters and re-tuning for every system update. An alternative is retrieval-augmented generation (RAG) (Lewis et al., 2020; Izacard and Grave, 2020; Borgeaud et al., 2022), which conditions output on information from an external source such as a dense retriever (Section 3.1). RAG methods facilitate online updates, reduce hallucinations, and generally require fewer LLM parameters since knowledge is externalized (Mialon et al., 2023; Borgeaud et al., 2022; Izacard and Grave, 2020).
RAG has recently begun to be explored for recommendation, with the most common approach being to first use a retriever or RS to construct a candidate item set based on a user query or interaction history, and then prompt an encoder-decoder LLM to rerank the candidate set (Hou et al., 2023; Wei et al., 2024; Wang and Lim, 2023; Dai et al., 2023; Yang et al., 2022). For RAG-based explanation generation, Xie et al. (Xie et al., 2023) generate queries based on interaction history to retrieve item reviews which are used as context to generate an explanation of the recommendation. RAG is also emerging as a key paradigm in conversational recommendation (c.f. Sec 3.5): for example, RAG is used in (Friedman et al., 2023) to retrieve relevant user preference descriptions from a user “memory” module to guide dialogue, and by Kemper et al. (Kemper et al., 2024) to retrieve information from an item’s reviews to answer user questions.
3.4. LLM-based Feature Extraction
Conversely to how RS or retrievers are used in RAG to obtain inputs for LLMs (Section 3.3), LLMs can also be used to generate inputs for RS (Harte et al., 2023; He et al., 2022b; Rajput et al., 2024; Yuan et al., 2023; Mysore et al., 2023; Li et al., 2023c). For instance: LLM2-BERT4Rec (Harte et al., 2023) initializes BERT4Rec (Section 2.1.1) item embeddings of item texts; Query-SeqRec (He et al., 2022b) includes LLM query embeddings as inputs to a transformer-based recommender; and TIGER (Rajput et al., 2024) first uses an LLM to embed item text, then quantizes this embedding into a semantic ID, and finally trains a T5-based RS to generate new IDs given a user’s item ID history. Similarly, MINT (Mysore et al., 2023) and GPT4Rec (Li et al., 2023c) produce inputs for a dense retriever by prompting an LLM to generate a query given a user’s interaction history.
@@ 3.5. Conversational Recommendation
The recent advances in LLMs have made fully NL system-user dialogues a feasible and novel recommendation interface, bringing in a new stage of conversational recommendation (ConvRec) research. This direction studies the application of LLMs in multi-turn, multi-task, and mixed-initiative NL recommendation conversations (Friedman et al., 2023; Jannach et al., 2020), introducing dialogue history as a rich new form of interaction data. Specifically, ConvRec includes the study and integration of diverse conversational elements such as dialogue management, recommendation, explanation, QA, critiquing, and preference elicitation (Jannach et al., 2020; Lyu et al., 2021). While some research (He et al., 2023) approaches ConvRec with a monolithic LLM such as GPT4, other works rely on an LLM to facilitate NL dialogue and integrate calls to a recommender module which generates item recommendations based on dialogue or interaction history (Li et al., 2018; Kang et al., 2019; Chen et al., 2019b; Yang et al., 2022; Wang et al., 2022b; Austin et al., 2024; Handa et al., 2024). Further research advances ConvRec system architectures with multiple tool-augmented LLM modules, incorporating components for dialogue management, explanation generation, and retrieval (Friedman et al., 2023; Gao et al., 2023; Wang et al., 2023a; Kemper et al., 2024; Zeng et al., 2024; Joko et al., 2024).
@@ 4. Generative Multimodal Recommendation Systems
In recent years, users have come to expect richer interactions than simple text or image queries. For instance, they might provide a picture of a desired product along with a natural language modification (e.g., a dress like the one in the picture but in red). Additionally, users want to visualize recommendations to see how a product fits their use case, such as how a garment might look on them or how a piece of furniture might look in their room. These interactions require new RS that can discover unique attributes in each modality. In this section, we discuss RS that utilize multiple data modalities. In Sections 4.1-4.2 we discuss motivations and challenges to the design of multimodal RS. In Sections 4.3-4.4 we review contrastive and generative approaches to multimodal RS, respectively.
@@ 4.1. Why Multimodal Recommendation?
A third use case for multimodal understanding is in RS with complex outputs, such as virtual try-on features or intelligent multimodal conversational shopping assistants.
@@ 4.2. Challenges to Multimodal Recommendation
The development of multimodal RS faces several challenges. First, collecting data to train multimodal systems (e.g., image-text-image triplets) is significantly harder than for unimodal systems. As a result, annotations for some modalities may be incomplete (Rahate et al., 2022).
Second, combining different data modalities to improve recommendation results is not simple. For instance, existing contrastive learning approaches (Radford et al., 2021; Jia et al., 2021; Li et al., 2021, 2022a) map each data modality to a common latent space in which all modalities are approximately aligned. However, such approaches often capture information that is shared across modalities (e.g., text describing visual attributes), but they overlook complementary aspects that could benefit recommendations (e.g., text describing non visual attributes) (Guo et al., 2019). In general we would like the modalities to compensate for one another and result in a more complete joint representation. While fusion-based approaches (Li et al., 2021, 2022a) do learn a joint multimodal representation, ensuring the alignment of information that is shared and leaving some flexibility to capture complementary information across modalities remains a challenge. Third, learning multimodal models requires orders of magnitude more data than learning models for individual data modalities.
@@ 4.4. Generative Multimodal Recommendation
Diffusion models, explained in Section 2.4, are state-of-the-art models for image generation. While they can also be used for text generation, e.g., by using a discrete latent space with categorical transition probabilities (Austin et al., 2021), text encoders based on transformers or other sequence-to-sequence models are preferred in practice. As a consequence, multimodal models for both text and images, such as text-to-image generation models, combine text encoders with diffusion models for images. For instance, DALL-E (Ramesh et al., 2022) uses the CLIP embedding space as a starting point to generate novel images, and Stable Diffusion (Rombach et al., 2022) uses a UNet autoencoder separately pre-trained on a perceptual loss and a patch-based adversarial objective. Several works have built on and expanded diffusion models by increasing controllability of the generated results (Zhang et al., 2023d), consistency on the generated subjects identity (Ruiz et al., 2023), or for virtual try on (Zhu et al., 2023).
@@ 5.1. Evaluating for Offline Impact
The typical approach to evaluating a model involves understanding its accuracy in an offline setting, followed by live experiments.
@@ 5.1.1. Accuracy Metrics
The usual metrics used for discriminative tasks are recall@k, precision@k, NDCG@k, AUC, ROC, RMSE, MAE, etc. Many recent works on generative RS (e.g., (Kang et al., 2017, 2023; Rajput et al., 2024; Bao et al., 2023; He et al., 2023)) incorporate such metrics for discriminative tasks.
@@ 5.1.3. Benchmarks
Many existing benchmark datasets popular in discriminative recommender models, such as Movielens (Harper and Konstan, 2015), Amazon Reviews (He and McAuley, 2016), Yelp Challenge(yel, 2014), Last.fm (Schedl, 2016), and Book-Crossing (Ziegler et al., 2005), are still useful in generative recommender models, but only narrowly. Some recent ones, like ReDial (Li et al., 2018) and INSPIRED (Hayati et al., 2020), are useful datasets for conversational recommendations. (Deldjoo, 2024a; Deldjoo and Di Noia, 2024; Zhang et al., 2023a) propose benchmarks called cFairLLM and FaiRLLM, to evaluate consumer fairness in LLMs based on the sensitivity of pretrained LLMs to protected attributes in tailoring recommendations. We note that some benchmarks such as BigBench(bench authors, 2023) which are commonly used by the LLM community, have recommendations tasks. It will be specifically useful for the RS community to develop new benchmarks for tasks unlocked by Gen-RecSys models.
@@ 5.2. Online and Longitudinal Evaluations
Offline experiments may not capture an accurate picture because of the interdependence of the different models used in the system and other factors. So, A/B experiments help understand the model’s performance along several axes in real-world settings. Note that (Wang et al., 2023d) proposes a new paradigm of using simulation using agents to evaluate recommender models. In addition to the short-term impact on engagement/satisfaction, the platform owners will be interested in understanding the long-term impact. This can be measured using business metrics such as revenue and engagement (time spent, conversions). Several metrics could be used to capture the impact on users (daily/monthly active users, user sentiment, safety, harm).
@@ 5.3. Conversational Evaluation
BLEU and perplexity are useful for conversational evaluation but should be supplemented with task-specific metrics (e.g., recall) or objective-specific metrics (e.g., response diversity (Li et al., 2015)). Strong LLMs can act as judges, but human evaluation remains the gold standard. Toolkits like CRSLab (K. Zhou, 2021) simplify building and evaluating conversational models, but lack of labeled data in industrial use cases poses a challenge. Some studies use LLM-powered user simulations to generate data.
5.4. Evaluating for Societal Impact
Previous work has investigated categories of interest for societal impacts of traditional RS (Milano et al., 2020) and generative models (Weidinger et al., 2022; Bird et al., 2023) independently. In the context of RS literature, six categories of harms are found to be associated with RS: content, privacy violations and data misuse, threats to human autonomy and well-being, transparency and accountability, harmful social effects such as filter bubbles, polarisation, manipulability, and fairness. In addition, RS based on generative models can present new challenges (Weidinger et al., 2022; Bird et al., 2023):
  • LLMs use out-of-domain knowledge, introducing different sources of societal bias that are not easily captured by existing evaluation techniques (Shen et al., 2023; Deldjoo, 2024b, a).
  • The significant computational requirements of LLMs lead to heightened environmental impacts (Luccioni et al., 2023; Berthelot et al., 2024).
@@ 5.4. Evaluating for Societal Impact
  • Recommender systems powered by generative models may be susceptible to manipulation and could have unintended and unexpected consequences for users (Carroll et al., 2022; Kasirzadeh and Evans, 2023).
  • Generative recommendations can expose users to the potential pitfalls of hyper-personalization (Gabriel et al., 2024; Rillig and Kasirzadeh, 2024).
@@ 5.5. Holistic Evaluations
As mentioned above, thoroughly evaluating RS for offline metrics, online performance, and harm is highly non-trivial. Moreover, different stakeholders (e.g. platform owners and users) (Abdollahpouri et al., 2020; Sürer et al., 2018; Milano et al., 2021) may approach evaluation differently. The complexity of Gen-RecSys evaluation presents an opportunity for further research and specialized tools. Drawing inspiration from the HELM benchmark (Liang et al., 2022), a comprehensive evaluation framework tailored for Gen-RecSys would benefit the community.
@@ 6. Conclusions and Future Directions
  • Red-teaming – in addition to the standard evaluations, real-world generative RS will have to undergo red-teaming (i.e., adversarial attacks) (Deng et al., 2023; Wang et al., 2023e; Shu et al., 2024) before deployment to stress test the system for prompt injections, robustness, alignment verification, and other factors.
Despite being a short survey, this work has attempted to provide a foundational understanding of the rich landscape of generative models within recommendation systems. It extends the discussion beyond LLMs to a broad spectrum of generative models, exploring their applications across user-item interactions, textual data, and multimodal contexts. It highlights key evaluation challenges, addressing performance, fairness, privacy, and societal impact, thereby establishing a new benchmark for future research in the domain.